### v2

# Chiral order emergence driven by quenched disorder

Abstract
Quenched disorder is expected to destroy order, such as when a random field is applied in a 2-dimensional magnetic system. Even when order exists in the presence of disorder, it is usually only the survival of the order of the clean model. We present here a surprising phenomenon where an order, driven by quenched disorder, emerges, that has nothing in common with the order present in the clean model. It’s a new type of order by disorder that differs from the usual thermal or quantum one. The classical Heisenberg model on the $J_{1}-J_{3}$ kagome lattice is studied by parallel tempering Monte Carlo simulations, with magnetic site vacancies. After analyzing the effect of a few impurities on the ground state, favoring non-coplanar configurations, we show the emergence of a low-temperature chiral phase and the progressive destruction of the collinear $q=4$ Potts order, the only order present in the clean system.
In frustrated two-dimensional magnets with continuous spins, discrete symmetries can be broken and lead to finite temperature phase transitions: time-reversal or lattice symmetry breaking induce chirality, stripe orders or valence bond crystals[1, 2]. Quenched disorder is expected to destroy or weaken such emerging order. However, on the classical kagome spin model with first and third neighbor Heisenberg exchanges ($J_{1}J_{3}$), we show that site disorder not only breaks the discrete $K_{4}$ order of the pure system[3], but but also brings about the appearance of a completely new chiral order (see Fig. 1). This phenomenon could turn out to be quite general, and is the consequence of two pieces that had previously been discovered but not assembled: i) Disorder can be the origin of order, as in the well-known case of thermal and quantum fluctuations[4], ii) Holes in a spin lattice favor the least-collinear configurations[5, 6].
Figure 1: Phase diagram of the $J_{1}J_{3}$ model for $J_{1}=-1$ and $J_{3}=1$. $p$ is the rate of magnetic vacancies, $T$ is the temperature. In the blue region the chirality $\chi\neq 0$. In the green region the Potts parameter $\sigma\neq 0$. The dots and crosses indicate phase transitions.
In most cases the broken symmetries in a clean system can be determined from the set of ground states. When they are not all related by a symmetry of the Hamiltonian, we speak from accidental degeneracies, as in the case of classical spin liquids[7]. At finite temperature, entropic effects favor the subset of ground states with the highest density of low-energy excitations. This is called the thermal order by disorder (ObD) phenomena, which also occurs in its quantum version at zero temperature for finite spin values due to quantum fluctuations.
The kagome antiferromagnet is the paradigmatic model of ObD, embodied by the $S=1/2$ compound Herbertsmithite[8], where both thermal and quantum fluctuations select coplanar states among its extensive number of ground states[9, 10, 11]. Among the many other kagome compounds (Kapellasite[12], Volborthite[13], Haydeite[14, 15], Sr-Vesignieite[16, 16]…), Ba-Vesignieite BaCu3V2O8(OH)2[17, 18, 19, 20, 21] can be described by a classical spin model which also presents ObD, where we now study the effect of site dilution.
This $J_{1}J_{3}$ model has ferromagnetic Heisenberg interactions $J_{1}$ between nearest neighbors and significant antiferromagnetic ones $J_{3}$ between third neighbors[22] (see Fig. 2), with a Curie-Weiss temperature of about $-77K$:
$\mathcal{H}=J_{1}\sum_{\langle i,j\rangle}\eta_{i}\eta_{j}\mathbf{S}_{i}\cdot\mathbf{S}_{j}+J_{3}\sum_{\langle i,j\rangle_{3}}\eta_{i}\eta_{j}\mathbf{S}_{i}\cdot\mathbf{S}_{j},$ (1)
where $\mathbf{S}_{i}$ are 3d unit vectors, $\eta_{i}$ is equal to $1$ for a magnetic site and 00 for a non-magnetic site (vacancy). The $\eta_{i}$ are statistically independent (uncorrelated disorder). $p=1-\sum_{i}\eta_{i}/N$ is the dilution parameter, with $N$ the total number of sites. Site dilution is an example of quenched disorder. We recall that a phase transition in spin model with quenched disorder usually falls into one of the following two categories, depending on the order parameter and on the type of quenched disorder[23]: (i) random field disorder, which arises when the disorder lifts the degeneracy between different ordered states, explicitly breaking the symmetry. As a consequence of the Imry-Ma argument [24], this prevents long-range order in dimension $d\leq 2$, (ii) random $T_{c}$ disorder, which preserves the broken-symmetry phase, but introduces spatial variations in the critical temperature. According to the Aizenman-Wehr argument[25], a phase transition can still occur if the pure model satisfies $\nu d<2$, where $\nu$ is the correlation length exponent $\xi\sim|T-T_{c}|^{-\nu}$. A representative example of these two categories is the $J_{1}-J_{2}$ Ising model on the square lattice, where site dilution affects the low-temperature phases (which depend on the values of $J_{1}$ and $J_{2}$) differently. While it preserves the ferromagnetic phase (acting as a random $T_{c}$ disorder), it destroys the stripe phase (acting as a random field disorder), by inducing a local breaking of the $C_{4}$ rotational lattice symmetry[26, 23, 27].
Figure 2: Left: $J_{1}$ and $J_{3}$ interactions considered in the $J_{1}J_{3}$ model. Right: collinear $K_{4}$ order, with its unit cell represented by a dashed line.
Theoretical studies of the pure $J_{1}J_{3}$ model have revealed a complex phase diagram[3, 28]. For $J_{1}=0$, the lattice decouples into three independent square sublattices, each exhibiting either ferromagnetic ($J_{3}<0$) or antiferromagnetic ($J_{3}>0$) order, leading in a 3-sublattice (for $J_{3}<0$) or 6-sublattice order (for $J_{3}>0$) (see Fig. 3). Beyond the expected $O(3)$ global spin rotational symmetry, an additional accidental degeneracy exists. When $J_{3}>0$, this accidental degeneracy is not lifted by a small $J_{1}$ (whatever its sign). The ground state manifold is called the 3subAF phase (see Fig. 3). By setting $J_{3}=1$, the range of $J_{1}$ values where this phase survives is quite extended[3]: $-1.24\simeq-\frac{1+\sqrt{5}}{4}<J_{1}<1$. We fix the value of $J_{3}=1$ from now on. However, the aforementioned ObD lifts this degeneracy at low temperature, stabilizing a discrete collinear $K_{4}$ (or equivalently $q=4$ Potts) order. The emergence of such a discrete order parameter in two dimensions via ObD has been observed in other models, such as the XY or Heisenberg $J_{1}-J_{2}$ square lattice[29, 30], where two columnar states are selected from the ground state manifold, leading to a $\mathbb{Z}_{2}$ order. A similar phenomenon was recently reported in the $J_{1}-J_{2}-J_{3}$ square lattice[31] with an analogous $K_{4}$ symmetry breaking.
Figure 3: Spin orientation in the 3subAF phase on the kagome lattice. Circles indicates the three sublattices $A$, $B$ and $C$, related by the $J_{3}$ interaction (red, blue and green links). On each of the sublattices, spins have two opposite orientations, forming three independently oriented Néel orders. When all spins are collinear, we get the order of Fig. 2.
Inevitably, perturbations - such as spin anisotropies - are present in realistic systems and can lift the ground state degeneracy, potentially favoring phases with alternative broken symmetries. However, the $K_{4}$ phase of the J1J3HK can persist at moderate temperatures if its entropy advantage compensates the energy cost. We focus here on a specific perturbation: quenched disorder, which can induce unexpected effects[32, 33, 5], that are fundamentally different from uniform perturbations of the Hamiltonian. In the J1J3HK model, it occurs that a discrete order parameter, chirality, emerges from site disorder, because it favors non-collinear configurations. The time-reversal symmetry is broken at low temperature, leading to a finite temperature phase transition. This reveals an unexpected result: a chiral ordered phase arising from quenched disorder, which is fundamentally different from the $K_{4}$ phase stabilized by entropic effects. Importantly, this phenomenon differs from a mere symmetry reduction that selectively eliminates certain ground states[27].
We present the phenomena selecting non-collinear configurations and then the numerical results evidencing the transition. Each configuration in the 3subAF phase is fully characterized by 3 spin orientations $(\mathbf{S}_{A},\mathbf{S}_{B},\mathbf{S}_{C})$ on a reference triangle $ABC$ of the lattice, as it allows to propagate these orientations and their opposites on the whole lattice (Fig. 3). Global spin rotation are symmetries of the model, thus ground states related by such transformations are considered as equivalent (this continuous symmetry is only broken at $T=0$ according to the Mermin Wagner theorem). The equivalence class of a ground state is is thus specified by a couple $(\bm{\sigma},\chi)$ of a three dimensional real vector $\bm{\sigma}$ and a chirality $\chi=\pm 1$ defined on a triangle $ABC$ as:
$\bm{\sigma}_{ABC}=\begin{pmatrix}\mathbf{S}_{B}\cdot\mathbf{S}_{C}\\ \mathbf{S}_{C}\cdot\mathbf{S}_{A}\\ \mathbf{S}_{A}\cdot\mathbf{S}_{B}\end{pmatrix},\quad\chi_{ABC}=\mathbf{S}_{A}\cdot(\mathbf{S}_{B}\land\mathbf{S}_{C}).$ (2)
Among these classes, 4 have collinear spins ±$\pm\mathbf{S}_{A}=\pm\mathbf{S}_{B}=\pm\mathbf{S}_{C}$ resulting in 4 possible values for $\bf\sigma$ and $\chi=0$ (represented on Fig. 2 right) and 2 others have orthogonal spins pointing to the vertices of an octahedron ($\bf\sigma=0$ and $\chi=\pm 1$). The configurations with orthogonal spins are called octahedral. These two chiral spin states (among all other chiral ones) have the particularity to not break any lattice symmetry[34].
At low temperature and for $p=0$, the 4 collinear classes of the 3subAF states are favored through thermal ObD[3]. It induces a discrete symmetry breaking related to the choice of one of the four possible classes (emergent $q=4$ Potts, or $K_{4}$ order parameter). The order parameter for this order is denoted $\Sigma$ and is defined in [3], while the chiral order parameter is $\chi$: the average of $\chi_{ABC}$ over the lattice. Finite-size analysis have shown that the critical exponents are compatible with the $4$ state Potts model universality class[3].
What are now the effect of a non-zero (small) $p$ in the $J_{1}J_{3}$ And first of all, does the $K_{4}$ order survive such disorder ? A random field coupled to a discrete order parameter is expected to destroy it in $d\leq 2$. In our case, although the disorder at hand is not a random field but site disorder, it still generates an effective random field, because both the $K_{4}$ order and the site vacancies break the translational lattice symmetries[23]. To illustrate this, we evaluate the effect of a small number of vacancies on the collinear spin configuration of Fig. 2 in the approximation where the state does not reorder to minimize its energy. It happens that a hole on a hexagon of green sites has the same energetic effect as a hole on a triangle of blue sites. However, as soon as two neighboring sites are empty, the energy difference with respect to the fully occupied state depends on the hole position: $8J_{3}-J_{1}$ for two sites of different colors versus $8J_{3}+J_{1}$ for two identical-color sites. Site disorder thus affects the 4 translationally equivalent collinear orders differently, similar to what a random field does on otherwise equivalent magnetizations. Thus, the $K_{4}$ order is expected to be killed by an infinitesimal $p>0$. We will see that this seductive scenario is challenged by the numerical simulations.
Figure 4: Ground state spin-spin correlation $\mathbf{S}_{i}\cdot\mathbf{S}_{0}$ between sites $i$ of the $A$ and $C$ sublattices and the central site 00 belonging to the $B$ sublattice. The exchanges of the $J_{1}$ and $J_{3}$ links emanating from site 00 tend to zero. Correlations between $B$ sites are not shown (they are strongly antiferromagnetic and near ±$\pm 1$.
Whatever, the competition between the 4 collinear states becomes completely irrelevant at very low temperature, as the energy minimization near vacancies kills any spin collinearity and allows the emergence of a new discrete order parameter, related to the chirality sign. Unlike $K_{4}$, chirality does not break any lattice symmetry and does not suffer from any random field effect, fully allowing a phase transition in the presence of site disorder. To illustrate this, we begin by considering a single vacancy on a $B$-site of the kagome lattice (Fig. 3). For $J_{1}=0$, the presence or absence of this spin has strictly no effect on the ground state manifold, but switching on $J_{1}$ lifts the degeneracy. The $A$ and $C$ spins of the four first neighbors of the impurity experience an apparent magnetic field of length $J_{1}$, in the direction opposite to the missing spin[35]. In a Néel antiferromagnet, parallel and perpendicular magnetic susceptibilities differ: $\mathcal{S}_{\parallel}=0$ while $\mathcal{S}_{\perp}>0$: the spins tend to orient perpendicularly to the magnetic field and to minimize their energy by a canting that decreases when moving away from the impurity. As a consequence, the $A$ and $C$ Néel ordered sub-lattices orient quasi-perpendicularly to the $B$ spins, with angles depending on the distance to the impurity (see correlations on Fig. 4) and alternating signs, similarly to Friedel oscillations already evidenced in other continuous spin models with single vacancies[36]. This phenomena is similar to the selection of the least-collinear order[5, 6], also observed on the $J_{1}-J_{2}$ square model[27, 37, 38], explained by an effective positive biquadratic interaction $(\mathbf{S}_{i}\cdot\mathbf{S}_{j})^{2}$ between neighboring spins. However, with a unique impurity, the $A$ and $C$ quasi-Néel orders remain mutually independent and we still expect them to align collinearly at inifinitesimal temperatures, a partial survival of the initial thermal ObD. With one impurity, we have only two perpendicular spin directions, which is not enough to specify a chirality. On the $J_{1}-J_{2}$ square lattice, the only 2 intricated Néel order prevent the apparition of chirality[38]. Here, with 3 sublattices and with vacancies on all types of sites, the 3 Néel orders tend to orient mutually perpendicularly, allowing to determine a local spin triedra with definite chirality.
We performed extensive Monte Carlo simulations for site dilution $0\leq p\leq 0.2$ and linear system size $L$ from $24$ to $128$ (the number of sites is $N=3L^{2}$). To minimize relaxation times, which can be significant at low temperatures, we employed a parallel tempering method combined with a local update provided by the heat bath algorithm [39]. Our numerical study focuses on the exchange parameters $J_{1}=-1$ and $J_{3}=1$ (we also investigated $J_{1}=-1.2$, resulting in a similar phase diagram[40]). For each $L$, simulations were performed with a minimum average of 10 disorder realizations.
The phase diagram $T-p$ is displayed in Fig. 1: the red dot at $p=0$ corresponds to the already known phase transition of the pure model, belonging to the universality class of the 4-state Potts model[3] (with critical exponents verifying $\alpha/\nu=1$, $\gamma/\nu=7/4$, and $21/\nu=3/$ 2), as clearly shown by finite-size scaling analysis (FSSA). We use the following relations for the maxima of the specific heat $C_{V}$ and of the (connected) susceptibility $S_{X}$ of the order parameter $X$ (with $X=\Sigma$ or $\chi$) to determine the nature of the phase transitions[41]:
$\begin{split}C_{V}^{\rm max}(L)&=\frac{\overline{\langle\mathcal{H}^{2}\rangle}-\overline{\langle\mathcal{H}\rangle^{2}}}{NT^{2}}\simeq aL^{\alpha/\nu}+b,\\ \mathcal{S}_{X}^{\rm max}(L)&=\frac{\overline{\langle X^{2}\rangle}-\overline{\langle X\rangle^{2}}}{NT}\simeq cL^{\gamma/\nu}+d,\\ T_{c}^{Y}(L)&\simeq eL^{-1/\nu}+T_{c}(\infty),\end{split}$ (3)
where brackets denote thermal averages and overlines are averages over disorder realizations. $T_{c}^{Y}$ is the temperature of the maximum of $Y$, where $Y$ is $C_{V}$, $\Sigma$ or $\chi$.
Figure 5: (left) Maxima of heat capacity $C_{V}^{\rm max}$ versus the linear lattice size $L$ and (right) $T_{c}^{C_{V}}(L)$ versus $L^{-3/2}$, for several dilution rates $p$.
Figure 6: (left) Order parameter $\Sigma$ (increasing and decreasing with $T$) and $\chi$ (immediately decreasing) versus $T$ for different dilutions and $L=96$. (right) Maximum of the susceptibilty $S_{\Sigma}^{\rm max}$ versus $L^{1.75}$. The dotted lines correspond to linear fits.
We now discuss the phase diagram for $p\neq 0$. For low impurity rate $p<0.05$, the transition temperature between the paramagnetic phase and the $K_{4}$ order decreases rapidly with increasing $p$. We observe a linear dependence of $C_{V}^{\rm max}(L)$ (Fig. 5 left), but the amplitude $a$ decreases rapidly with $p$. The evolution of $T_{c}^{C_{V}}(L)$ (Fig. 5 right) confirms that $\nu=2/3$. Lastly, the maximum of the $K_{4}$ susceptibilty $\mathcal{S}_{\Sigma}^{\rm max}$ increases as $L^{7/4}$ with a slope decreasing with $p$ (Fig. 6 right). Thus FSSA clearly indicates that there is a phase transition in the same universality class as the pure model, i.e as the 4-state Potts model. This result seems to invalidate the argument that site dilution leads to the existence of local random fields, but we want to draw attention on the fact that other studies report transitions in presence of random fields, arguing that their lattice-symmetry breaking effect is very small[26].
At lower temperatures, a second phase transition appears, where we observe simultaneously a rapid decrease of $\Sigma$ ($K_{4}$ order) and a rapid increase of $\chi$ (chiral order) as $T$ decreases (see Fig. 6 left). In the presence of site disorder, non-collinear configurations are energetically favored, overcoming the entropic advantage of the $K_{4}$ order at very low temperatures. The thermodynamic signature of this transition is the appearance of a small secondary maximum in $C_{V}$ at $T_{c}\simeq 4p$ for $p<0.05$, which increases weakly with $L$ (e.g., $C_{V}^{\rm max}(28)=1.15$, $C_{V}^{\rm max}(128)=1.20$). This prevents the determination of a reliable scaling exponent with FSSA, that however remains compatible with the scaling of the dilute Ising spin model[42]. $S_{\chi}$ appears to be weakly dependent on the system size $L$, while $S_{\Sigma}$ follows a scaling behavior with an exponent slightly larger than 1. To obtain stable results in this regime, a disorder average over $20$ realizations is required. Results are also also limited to $L\leq 80$, due to thermalization difficulties.
For $0.05\leq p\leq 0.1$, $\Sigma$ collapses to zero at all temperatures as $L$ increases. This behavior becomes more pronounced as $p$ increases[40]. In the thermodynamic limit, we expect the $K_{4}$ order to vanish, which may indicate that site dilution finally acts as a random field and kills the transition for $p>0.05$. At low temperatures, the chiral order is present below a critical temperature that increases with the dilution, but saturates at $p\simeq 0.1$ (see Fig. 1). For $0.1\leq p\leq 0.2$, the $K_{4}$ order is clearly absent, and the chiral order weakens as $p$ increases with a progressive decrease of the critical temperature. Nevertheless, one observes a significant versatility of behaviors: for $p=[0.14,0.16]$, $C_{V}$ displays a small but noticeable divergence with the system size[40]. Moreover, the maximum of $S_{\chi}$ still diverges with $L$, even though its value decreases due to the site dilution, which isolates some of the complete triangles where $\chi$ is defined. FSSA unveils that this maximum scales as $L^{1.7}$ which is compatible with the Ising universality class.
For $p=0.2$, the chiral order parameter collapses with $L$, which can be interpreted as the absence of a phase transition due to disorder overcoming some percolation threshold. On the kagome lattice, triangles form a honeycomb lattice and have a probability of $(1-p)^{3}$ to be fully occupied by magnetic sites. Equating in a crude approximation this value with the site percolation threshold of the honeycomb lattice (equal to 0.697[43]) gives a critical dilution $p_{c}=0.11$, but this has to be taken with care as further neighbor interactions complicate this view in term of percolation.
In summary, we have shown on the example of the $J_{1}J_{3}$ model with magnetic site vacancies that quenched disorder can lead to the emergence of a specific order (here chiral), that was forbidden in the clean phase (with collinear $K_{4}$ order). This phenomenon can be designated as the order by quench disorder mechanism that differs from the usual thermal or quantum order by disorder. Although this expression has been previously employed[6], our proposal is distinct in that it involves the breaking of time-reversal symmetry in the absence of a magnetic field and in contrast to the pure model. A similar birth of chirality is expected for other models such as the triangular or honeycomb lattices with stripe orders[44, 34], due to the lattice decoupling in more than two sublattices (similar to what occurs in the $J_{1}J_{3}$ model). The exploration of classical spin liquids[7] represents another direction for further research on order from quenched disorder.
I Acknowledgments
L.M thanks Eric Andrade for invaluable discussions on the effect of disorder, and the IIT Madras for its hospitality that made these discussions possible.